OpenStudy (anonymous):

Find the area of a kite with diagonals measuring 6 cm and 18 cm.

OpenStudy (campbell_st):

\[A = \frac{1}{2}\times x \times y\] x and y are the diagonal lengths... hope this helps

OpenStudy (anonymous):

so its 54?

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

don't forget the units...cm^2 and your answer is correct...

Directrix (directrix):

The area of a kite: To find the area of a kite, multiply the lengths of the two diagonals and divide by 2 (same as multiplying by 1/2): (1/2) times 6 times 18 @Foreverisabell >> so its 54? YES and don't forget @campbell_st 's reminder